Circular No. 7777

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2001 X2 (SCOTTI) Additional astrometry and the following extremely uncertain orbital

elements appear on MPEC 2001-Y14:

T = 2002 Aug. 10.311 TT  ω = 7.^{o}808
e = 0.47686  Ω = 183.124 }2000.0
q = 1.95226 AU  i = 2.334
a = 3.73181 AU n^{o} = 0.136718 P = 7.21 years

SUPERNOVA 2001ig IN NGC 7424 S. Ryder and K. Kranz, Anglo-Australian Observatory; E. Sadler, Uni-

versity of Sydney; and R. Subrahmanyan, Australia Telescope National Facility, write: “We have searched for the radio counterpart to SN 2001ig with the Australia Telescope Compact Array on Dec. 15.3–15.6 UT. A source was detected in an 8640-MHz image at α = 22^{h}57^{m}30^{s}.74, δ = −41^{o}02′26′′.1 (equinox 2000.0) with a flux of 2.2 ± 0.3 mJy. No source was detected at 4790 MHz above a level of 0.9 mJy, implying a spectral index > 1.5. The source was also observed on Dec. 18.1–18.2, by which time the flux at 8640 MHz had doubled, and a marginal detection at 4790 MHz of 0.9 mJy was obtained. Continued monitoring at these and other frequencies is planned in coming weeks.”

DK LACERTAE A. Henden, USRA and U.S. Naval Observatory (USNO); and E. Free-

land and K. Honeycutt, Indiana University, write: “Regular monitoring since 1990 of DK Lac (N Lac 1950) with Indiana’s RoboScope shows this cataclysmic variable to have been stable near V = 16.8 until a slow decline began during Sept. 2000. By Jan. 2001, the system had largely faded below the RoboScope detection magnitude limit of 17.5–18.5. Exposures on Dec. 15 UT with the WIYN 0.91-m telescope at Kitt Peak, and on Dec. 17 with the 1.0-m telescope at USNO/Flagstaff, show the system near V = 19.4 with B −V = −0.2. There is a red companion (mag 19.5) located 5′′ to the east. VY Scl-type low states such as these provide an opportunity to study the white dwarf while the accretion luminosity is reduced or absent.”

2001 December 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green